OpenStudy (mwhit2101):

SIXTH GRADE MATH HELP MEDAL PLUS FAN! The graph of which function passes through the point (3, 24)? a. y=2x^2+6 b. y=-2x^2+6 c. y=6x^2+2 d. y=2x^2-6

OpenStudy (muzzack):

well first put in the firsr two coodinates in x and y in each answer

OpenStudy (muzzack):

a) 24 = 2(3)^2+6 try it

OpenStudy (mwhit2101):

Ok...hold on

OpenStudy (muzzack):

b) 24=-2(3)^2+6

OpenStudy (mwhit2101):

How do I know when it's passing through? I'm more confused on that part

OpenStudy (muzzack):

c. 24=6(3)^2+2 d. 24=2(3)^2-6

OpenStudy (anonymous):

put in the first two coodinates in x and y in each answer

OpenStudy (mwhit2101):

How do I know when it's passing through? I'm more confused on that part

OpenStudy (mwhit2101):

@Muzzack

OpenStudy (anonymous):

DID U TRY THE EQUATION that he said ?

OpenStudy (mwhit2101):

OMG IM CONFUSED HOW DO I KNOW WHEN IT IS PASSING THROUGH I DONT GETI T

OpenStudy (mwhit2101):

nvm i figured it out without yall. bye.